OpenStudy (anonymous):

binomial probabilty According to the Website Anwers.com, 10-13% of Americans are left-handed. Assume that the percentage is 11%. If we select 9 people at random, find the probability that the number who are left handed is a, at least 2

OpenStudy (anonymous):

at least two means 2, or 3, or 4, or 5, or 6 or 7 or 8 or 9 which is way too many probabilities to compute can you think of a much faster way?

OpenStudy (anonymous):

i was thinking that i can find at most 1 and then subtract my answer from 100% @satellite73

OpenStudy (anonymous):

would i be right that way?

OpenStudy (anonymous):

not "at most one" but rather "exactly 1" also "exactly 0" find those two, add them up, and subtract the result from one to get the answer to "at least 2"

OpenStudy (anonymous):

exactly none is the easiest one do you know what it is?

OpenStudy (anonymous):

noo

OpenStudy (anonymous):

ok so we go slow 11% are left handed, what is the percent that is not left handed?

OpenStudy (anonymous):

.11

OpenStudy (anonymous):

100-11 =89

OpenStudy (anonymous):

\(11\%=.11\) yes

OpenStudy (anonymous):

and so \(89\%=.89\) for not left handed the probability that 9 in a row are not left handed is therefore \[P(x=0)=(.89)^9\] which you will need a calculator to compute

OpenStudy (anonymous):

i had the overall sing the binomial formula to be 0.3504

OpenStudy (anonymous):

we are not done we still need \[P(x=1)\]

OpenStudy (anonymous):

i had for the second 0.3897

OpenStudy (anonymous):

so far all we have computed is the probability that none are left handed probability that one is left handed is \[P(x=1)=9\times (.11)\times (.89)^8\]

OpenStudy (anonymous):

for that one i get \(.433\)

OpenStudy (anonymous):

for a final answer of \[1-.433-.3504\]

OpenStudy (anonymous):

oh ok..so i need not convert to any percentage right

OpenStudy (anonymous):

no you just need to subtract to get the probability if you want for some reason to convert to a percent, move the decimal over two places

OpenStudy (anonymous):

oh ok..

OpenStudy (anonymous):

u the best

OpenStudy (anonymous):

yw